b) Show that (1+i) is a zero of F(x) = 2x^5 - 9x^4 +12x^3 -4x^2 - 8x +4 c) Find all of the ZEROES of F(x)
Answer #1

b) Give F(x) = 2005 _ 9 oc + 12x²_40c²-83ct4 - Now Hi is a sero of the fox) if on substituting = Hi in f(x) it gives f (ti)=(© . Since Complex Heros are occurs in Conjugate bair therefore x=1-1 is also the seas of food. Now X= Hii and xal-i 26-1-1=0Now find other 3 years using by solving for 226?- Sz 2-2x+2=0 = By guess we found a = 1/2 is the solution of the above equati